# Efficient Tensor Completion Algorithms for Highly Oscillatory Operators

### v2

@@ 5.3 Full Tensor-Based Butterfly Completion
The overall workflow and cost of the proposed tensor completion algorithms for butterfly matrices are as follow. Assuming a butterfly rank $r$ and given $|\Omega|=\mathcal{O}(nr\log n)$ (which is the empirical estimate of the minimal number entries required by the butterfly completion), we first generate the initial guess by low-rank completion with a constant number of iterations and rank estimate $R$, followed by the low-rank to butterfly conversion (Algorithm 4). Unless otherwise stated, we use $R=r$. Hence the initial guess generation requires a total of $\mathcal{O}(nr^{4}\log n)$ operations. Next, we run the butterfly completion (ALS-based Algorithm 1 or ADAM-based Algorithm 3) for at most $t_{\max}$ iterations. Here $t_{\max}$ is also set to a constant. Each iteration requires $\mathcal{O}(|\Omega|r^{4}\log^{2}n+nr^{6}\log n)=\mathcal{O}(nr^{5}\log^{3}n)$ operations for ALS and $O(|\Omega|r^{2}\log^{2}n+nr^{2}\log n)=O(nr^{3}\log^{3}n)$ operations for ADAM. All of the above algorithms are implemented in Python and are open source [^1] https://github.com/navjo2323/butterfly-matrix-completion.
@@ 6 Numerical Experiments
This section provides several numerical examples to demonstrate the convergence, accuracy and efficiency of the proposed matrix completion algorithms, i.e., Algorithm 1 (ALS-based) and Algorithm 3 (ADAM-based), when applied to incomplete data from high-frequency Green’s function for wave equations (Section 6.1), inverse scattering applications (Section 6.2) and Radon transforms (Section 6.3). As the baseline algorithms, we use the ALS-based low-rank matrix completion algorithm (Algorithm 1 with $L=0$) and the ALS-based QTT tensor completion algorithm (Algorithm 5). A convergence tolerance of $\epsilon=10^{-3}$ is used for all the algorithms. All the proposed and baseline algorithms are implemented in Python and executed using one CPU processor of the Perlmutter machine at NERSC in Lawrence Berkeley National Laboratory, which has AMD EPYC 7763 processors and 128GB of 2133MHz DDR4 memory.
@@ 6.1 Green’s function for Helmholtz equation
First, we consider an example of $n=4096$ and compare the reconstruction quality of the butterfly completion algorithm (Algorithm 1) and low-rank matrix completion algorithm. The real part of the full matrix $\bm{T}$ and the observed submatrix $\bm{T}_{\Omega}$ are shown in Fig. 3(a) and (b), where the number of observed entries is empirically set to $|\Omega|=66n\log n=3244032$. The low-rank completion algorithm uses rank estimate $r=400$, max iterations $t_{max}=20$, and the reconstructed full matrix $\bm{X}$ is shown in Fig. 3(c). The butterfly completion algorithm (Algorithm 1) uses rank estimate $r=11$, max iterations $t_{max}=20$ and use the initial guess with $R=11$. The reconstructed full matrix $\bm{X}$ is shown in Fig. 3(d). As one can see, the reconstruction quality of Algorithm 1 is significantly better than that of the low-rank completion algorithm, given the same set of observed entries.
@@
Next, we examine the convergence of the completion algorithms. We first consider an example of $n=16384$ ($L=12$) using the low-rank matrix completion with rank $r=200$ or $400$, the ALS-based and gradient-based butterfly completion Algorithm 1 and Algorithm 3 with $r=11$ and $R=11$, and ALS-based QTT completion Algorithm 5 with rank $r=50$. The number of observed entries is empirically set to $|\Omega|=66n\log n$ for training and $|\Omega_{\mathrm{test}}|=11n\log n$ for testing. The convergence history for the training and test errors are shown in Fig. 4. We set max iterations $t_{max}=20$ for all the ALS-based algorithms and $t_{max}=100$ for the gradient-based algorithm. Note that the low-rank and QTT completion algorithms do not converge for the rank estimates used and will run out of memory for even larger rank estimates. In contrast, the butterfly completion algorithms converge with a small rank $r=11$. Specially, the ALS-based butterfly algorithm converges in 9 iterations, while the gradient-based butterfly algorithm takes more than 100 iterations to converge. Next, an even larger problem is tested as $n=65536$ ($L=14$). We only test the low-rank matrix completion algorithms with $r=200$ or $400$ and Algorithm 1 with $r=11$. As can be seen from Fig. 5, the low-rank matrix completion algorithms diverge but the butterfly algorithm converges in 16 iterations.
@@
Figure 4: The convergence of a Green’s function example with $n=16384$, $c=4$, $L=12$ using the low-rank matrix completion with rank $r=200,400$, butterfly completion Algorithm 1 and Algorithm 3 with rank $r=11$, and QTT completion Algorithm 5 with rank $r=50$. (a) Training error with $|\Omega|=66n\log n$. (b) Test error with $|\Omega_{\rm test}|=11n\log n$.
@@
Figure 5: The convergence of a Green’s function example with $n=65536$, $c=4$, $L=14$ using the low-rank matrix completion with rank $r=200$, and butterfly completion Algorithm 1 with rank $r=11$. (a) Training error with $|\Omega|=66n\log n$. (b) Test error with $|\Omega_{\rm test}|=11n\log n$.
@@
Next, we investigate the computational complexity of these completion algorithms with the minimum number of nonzeros it took for the algorithms to converge. As $L\in\{8,\dots,14\}$, we set the number of observed entries to $|\Omega|=10rn$ for the low-rank matrix completion algorithm, $|\Omega|=6n\log n$ for the butterfly completion algorithms, and $|\Omega|=3rn$ for the QTT algorithm. The rank estimates are set to the empirically smallest numbers such that the completion algorithms converge for a given $n$. More specifically, we set $r=43,100,253$ for the low-rank algorithm, $r=11$ for the butterfly algorithms and $r=30,54,96$ for the QTT algorithm. Note that for $n=65536$ ($L=14$), both low-rank matrix completion and QTT completion algorithms run out of memory. We plot the CPU time per iteration for each algorithm in Fig. 6(a) and the number of observed entries $|\Omega|$ in Fig. 6(b). Note that the time per iteration for ALS-based and ADAM-based butterfly algorithms almost overlap with each other. The butterfly algorithms emprically achieve the lowest total asymptotic complexity, $\mathcal{O}(n\log^{3}n)$, among all the completion algorithms.
@@
Figure 6: (a) Computational time per iteration and (b) $|\Omega|$ for a Green’s function example with $c=4$, L=814L=8-14 using the low-rank matrix completion, butterfly completion Algorithm 1 and Algorithm 3, and QTT completion Algorithm 5. Butterfly completion: $|\Omega|=6n\log n$, $r=11$. Low-rank completion: $|\Omega|=10rn$, $43,100,253$. Note that in theory low-rank representation requires $r=\mathcal{O}(n)$ for sufficiently large $n$, these numerical tests exhibit $r=\mathcal{O}(\sqrt{n})$ up to $n=16384$. QTT completion: $|\Omega|=3rn$, $r=30,54,96$. Note that in theory QTT representation requires $r=\mathcal{O}(\sqrt{n})$ for sufficiently large $n$, these numerical tests exhibit $r=\mathcal{O}(n^{0.3})$ up to $n=65536$.
@@ 6.2 Inverse wave scattering
To generate the Green’s function matrix $\bm{T}$, we leverage the simulation package WAVEFORM [76] to simulate the responses $p(\mathbf{x}^{o}_{j})$, $j=1,2,\ldots,n$ for each excitation function $f({\mathbf{x}})=\delta(\mathbf{x}-\mathbf{x}^{s}_{i})$. The frequency is set to $\omega=500\pi$ and $n$ is set to 1024 (i.e., $c=4$, $L=8$). Once the matrix is fully generated, we assume only $|\Omega|$ entries of the matrix are observed and our goal is to fully recover $\bm{T}$. For the transmission problem, we empirically set $|\Omega|=27n\log n$, $r=9$ for the butterfly completion Algorithm 1 and $r=120$ for the low-rank completion. For the reflection problem, we empirically set $|\Omega|=48n\log n$, $r=12$ for the butterfly completion Algorithm 1 and $r=120$ for the low-rank completion. The reconstructed matrices are plotted in the last two columns of Fig. 7. Clearly, the butterfly completion algorithm achieves significantly better reconstruction quality than the low-rank completion algorithm.
@@ 6.3 Radon transform
First, we consider an example of $n=1024$ and compare the reconstruction quality of the butterfly completion algorithm (Algorithm 1) and low-rank matrix completion algorithm. The real part of the full matrix $\bm{T}$ and the observed submatrix $\bm{T}_{\Omega}$ are shown in Fig. 8(a) and (b), where the number of observed entries is empirically set to $|\Omega|=30n\log n=307200$. The low-rank matrix completion algorithm uses rank estimate $r=400$, max iterations $t_{8}{max}=20$, and the reconstructed full matrix $\bm{X}$ is shown in Fig. 8(c). The butterfly completion algorithm (Algorithm 1) uses rank estimate $r=10$, max iterations $t_{max}=20$ and the initial guess with $R=10$. The reconstructed full matrix $\bm{X}$ is shown in Fig. 8(d). As one can see, the reconstruction quality of Algorithm 1 is significantly better than that of the low-rank completion algorithm, given the same set of observed entries. This is no surprise as the low-rank completion algorithm requires $r=\mathcal{O}(n)$ and $|\Omega|=\mathcal{O}(n^{2})$ for a good reconstruction.
@@
Next, we examine the convergence of the completion algorithms. We first consider an example of $n=4096$ ($L=10$) using the low-rank matrix completion with rank $r=100$ or $200$, and the ALS-based and gradient-based butterfly completion Algorithm 1 and Algorithm 3 with rank $r=12$. Note that the QTT completion algorithm is not considered as the full QTT rank will yield the algorithm prohibitively expensive. The number of observed entries is empirically set to $|\Omega|=84n\log n$ for training and $|\Omega_{\mathrm{test}}|=12n\log n$ for testing. The convergence history for the training and test errors are shown in Fig. 9. We set max iterations $t_{max}=20$ for all the ALS-based algorithms and $t_{max}=60$ for the gradient-based algorithm. As expected, the low-rank matrix completion algorithms do not converge. In contrast, the butterfly completion algorithms converge with a small rank $r=12$. Specially, the ALS-based butterfly algorithm converges in 16 iterations, while the gradient-based butterfly algorithm takes more than 60 iterations to converge.
@@
Figure 9: The convergence of a Radon transform example with $n=4096$, $c=4$, $L=10$ using the low-rank matrix completion with rank $r=100,400$, and butterfly completion Algorithm 1 and Algorithm 3 with rank $r=12$. (a) Training error with $|\Omega|=84n\log n$. (b) Test error with $|\Omega_{\rm test}|=12n\log n$.
@@
Figure 10: (a) Computational time per iteration and (b) $|\Omega|$ for the Radon transform example with $c=4$, L=614L=6-14 using the low-rank matrix completion, and butterfly completion Algorithm 1 and Algorithm 3. Butterfly completion: $|\Omega|=6n\log n$, $r=10$. Low-rank completion: $|\Omega|=n^{2}$ (requires all the entries to be decomposed), $r=n$ (full rank).
@@
Next, we investigate the computational complexity of these completion algorithms for the radon transform input matrices. As $L\in\{6,\dots,14\}$, we set the number of observed entries to $|\Omega|=n^{2}$ for the low-rank algorithm as it needs all the entries to be observed to converge to sufficient accuracy, and $|\Omega|=6n\log n$ for the butterfly algorithms. The rank estimate is $r=12$ for the butterfly algorithms and $r=n$ for the low-rank algorithm. Note that the low-rank algorithm runs out of memory beyond $n=4096$. The memory error can be delayed a bit further to use larger $n$ by using matrix decomposition algorithm as the algorithm needs all entries as input, the computation with full matrices is memory bound. Furthermore, this suggests that low-rank matrix completion is not suitable for radon transform input matrices. We plot the CPU time per iteration for each algorithm in Fig. 10(a) and the number of observed entries $|\Omega|$ in Fig. 10(b). Note that the time per iteration for ALS-based and ADAM-based butterfly algorithms almost overlap with each other. Clearly, the butterfly algorithms achieve the quasi-linear asymptotic complexity as $\mathcal{O}(n\log^{3}n)$.

### v3

@@ 5.3 Full Tensor-Based Butterfly Completion
The overall workflow and cost of the proposed tensor completion algorithms for butterfly matrices are as follow. Assuming a butterfly rank $r$ and given $|\Omega|=\mathcal{O}(nr\log n)$ (which is the empirical estimate of the minimal number entries required by the butterfly completion), we first generate the initial guess by low-rank completion with a constant number of iterations and rank estimate $R$, followed by the low-rank to butterfly conversion (Algorithm 4). Unless otherwise stated, we use $R=r$. Hence the initial guess generation requires a total of $\mathcal{O}(nr^{4}\log n)$ operations. Next, we run the butterfly completion (ALS-based Algorithm 1, ADAM-based Algorithm 3, or ADF-based) for at most $t_{\max}$ iterations. Here $t_{\max}$ is also set to a constant. Each iteration requires $\mathcal{O}(|\Omega|r^{4}\log^{2}n+nr^{6}\log n)=\mathcal{O}(nr^{5}\log^{3}n)$ operations for ALS and $O(|\Omega|r^{2}\log^{2}n+nr^{2}\log n)=O(nr^{3}\log^{3}n)$ operations for ADAM and ADF. As a baseline, we also implement several QTT tensor completion algorithms that are ALS, ADAM or ALS-based. All of the above algorithms are implemented in Python and are available online  [^1] https://github.com/navjo2323/butterfly-matrix-completion.
A remark on sampling complexity. We remark that identification of the sampling complexity of various matrix and tensor completion algorithms is a still active research area. Ideally, the number of observed entries $|\Omega|$ should scale linearly with the number of degrees of freedom of the underlying matrix or tensor representation. While practical algorithms for completion have a long history of development, theoretical guarantees regarding their optimal sampling complexity have only recently been established. For low-rank matrix completion, earlier works can reach the optimal sampling complexity $|\Omega|=\mathcal{O}(rn)$ only for nuclear norm-based algorithms [recht2011simpler, mu2014square], but sub-optimal complexity for ALS-based [jain2015fast] algorithms. The optimal complexity for ALS-based algorithms has only been recently proved [kelner2023matrix]. For QTT (or TT) completion, a very recent work proves the optimal $|\Omega|=\mathcal{O}(r^{2}\log n)$ sampling complexity for Riemannian optimization-based completion algorithms [budzinskiy2023tensor]. As our paper represents the first completion algorithms for butterfly representation, we do not attempt to establish a theoretical bound for the sampling complexity, but rather focus on empirically demonstrating the effectiveness of the proposed butterfly completion algorithms for highly oscillatory matrices. We empirically show that the sampling complexity of $|\Omega|=\mathcal{O}(rn\log n)$ is sufficient for several numerical examples in Section 6 and leave its theoretical justification as a future work.
@@ 6 Numerical Experiments
This section provides several numerical examples to demonstrate the convergence, accuracy and efficiency of the proposed completion algorithms when applied to incomplete data from high-frequency Green’s function for wave equations (Section 6.1), inverse scattering applications (Section 6.2) and Radon transforms (Section 6.3). Specifically, the proposed algorithms are butterfly completion algorithms that are ALS-based (Algorithm 1), ADAM-based (Algorithm 3), and ADF-based (Algorithm 6). As for the baseline algorithms, we use the ALS-based low-rank matrix completion algorithm (Algorithm 1 with $L=0$) and the QTT tensor completion algorithms that are ALS-based (Algorithm 5), ADAM-based and ADF-based. A convergence tolerance of $\epsilon=10^{-3}$ is used for all the algorithms. All the proposed and baseline algorithms are implemented in Python and executed using one CPU processor of the Perlmutter machine at NERSC in Lawrence Berkeley National Laboratory, which has AMD EPYC 7763 processors and 128GB of 2133MHz DDR4 memory.
@@ 6.1 Green’s function for Helmholtz equation
First, we consider an example of $n=4096$ and compare the reconstruction quality of the butterfly completion algorithm (Butterfly-ALS of Algorithm 1) and low-rank matrix completion algorithm. The real part of the full matrix $\bm{T}$ and the observed submatrix $\bm{T}_{\Omega}$ are shown in Fig. 3(a) and (b), where the number of observed entries is empirically set to $|\Omega|=66n\log n=3244032$. The low-rank completion algorithm uses rank estimate $r=400$, max iterations $t_{max}=20$, and the reconstructed full matrix $\bm{X}$ is shown in Fig. 3(c). The butterfly completion algorithm (Algorithm 1) uses rank estimate $r=11$, max iterations $t_{max}=20$ and use the initial guess with $R=11$. The reconstructed full matrix $\bm{X}$ is shown in Fig. 3(d). As one can see, the reconstruction quality of Algorithm 1 is significantly better than that of the low-rank completion algorithm, given the same set of observed entries.
@@
Next, we examine the convergence of the completion algorithms using the real part of the Green’s function as $\bm{T}$. We first consider an example of $n=16384$ ($L=12$). As the ground truth, the matrix rank, butterfly rank and QTT rank with tolerance $\epsilon=10^{-3}$ are 366, 11 and 96, respectively. For all algorithms in this example, we set the rank estimates in the completion algorithms as the above true numerical ranks. Specifically, for low-rank completion, we use $r=366$ and $|\Omega|=10rn=59965440$; for the ALS-based (Algorithm 1), gradient-based (Algorithm 3) and ADF-based(Algorithm 6) butterfly completion, we use $r=11$, $R=11$ and $|\Omega|=6rn\log n=15138816$; for the ALS-based (Algorithm 5), gradient-based and ADF-based QTT completion, we set $r=96$, $R=11$ and $|\Omega|=16r^{2}\log n=2240000$. The convergence history for the training and test errors are shown in Fig. 4. We set max iterations $t_{max}=20$ for all algorithms. As we set $|\Omega|$ and $r$ to approximately the minimum numbers required by each completion algorithm, all algorithms show convergence. Specifically, gradient-based methods such as Butterfly-ADAM and QTT-ADAM show slowest convergence; ADF-based algorithms show slightly worse convergence behaviors than ALS-based algorithms. For low-rank matrix completion, we also plot the convergence history for $|\Omega|=4rn=23986176$ (labeled as “low-rank(low)”), which appears to be insufficient. For QTT-ADF, we also plot the convergence history for $|\Omega|=8r^{2}\log n=1120000$ (labeled as “QTT-ADF(low)”), which also turns to be an insufficient number for convergence. Moreover, although QTT completion requires the lowest $|\Omega|$, the cost per iteration (see Fig. 6 for more detail) is most expensive (e.g., at $L=12$, QTT-ADF is $10\times$ slower than Butterfly-ALS, and QTT-ALS is $50\times$ slower than Butterfly-ALS) and the test errors are less satisfactory than butterfly completion. Next, an even larger problem is tested as $n=65536$ ($L=14$). We only test the low-rank matrix completion algorithms with $r=200$ or $400$ and Butterfly-ALS/Butterfly-ADF with $r=11$. We skip Butterfly-ADAM due to its slow convergence and QTT-completion due to its high memory usage and computational cost attributed to large QTT ranks. We set $|\Omega|=69206016$ for both low-rank and butterfly completion algorithms. As can be seen from Fig. 5, the low-rank matrix completion algorithms diverge due to both insufficient rank estimate and $|\Omega|$ (as will be seen in Fig. 6, it requires a projected 4 hours per iteration for low-rank completion to converge). In contrast, the butterfly algorithm converges within 20 iterations. Again, the convergence of Butterfly-ADF is slightly worse than Butterfly-ALS.
@@
Figure 4: The convergence of a Green’s function example with $n=16384$, $c=4$, $L=12$ using the low-rank matrix completion with $r=366$, $|\Omega|=10rn=59965440$ (i.e., $|\Omega|/n^{2}=0.22$), and $|\Omega|=4rn=23986176$ (i.e., $|\Omega|/n^{2}=0.09$, labeled as “Low-rank(low)”), butterfly completion (ALS, ADAM, ADF) with $r=11$, $|\Omega|=6rn\log n=15138816$ (i.e., $|\Omega|/n^{2}=0.056$), and QTT completion (ALS, ADAM, ADF) with $r=96$, $|\Omega|=16r^{2}\log n=2240000$ (i.e., $|\Omega|/n^{2}=0.0083$). In addition, “QTT-ADF(low)” denotes ADF-based QTT completion with $|\Omega|=8r^{2}\log n=1120000$. (a) Training error. (b) Test error.
@@
Figure 5: The convergence of a Green’s function example with $n=65536$, $c=4$, $L=14$ using butterfly completion (ALS and ADF) with rank $r=11$. (a) Training error with $|\Omega|=6rn\log n=69206016$. (b) Test error with $|\Omega_{\rm test}|=rn\log n=11534336$. The results using low-rank matrix completion with rank $r=200,400$ and the same $|\Omega|$ as butterfly completion are also plotted.
@@
Next, we investigate the computational complexity of these completion algorithms with the minimum number of nonzeros it took for the algorithms to converge. As $L\in\{8,\dots,14\}$, we set the number of observed entries to $|\Omega|=10rn$ for the low-rank matrix completion algorithm, $|\Omega|=6n\log n$ for the butterfly completion algorithms, and $|\Omega|=16r^{2}\log n$ for the QTT algorithms. The rank estimates are set to the empirically smallest numbers such that the completion algorithms converge for a given $n$. More specifically, we set $r=57,134,366$ for the low-rank algorithm, $r=11$ for the butterfly algorithms and $r=30,54,96$ for the QTT algorithm. Note that for $n=65536$ ($L=14$), both low-rank matrix completion and QTT completion algorithms run out of memory or are too expensive to finish. We plot the CPU time per iteration for each algorithm in Fig. 6(a) and the number of observed entries $|\Omega|$ in Fig. 6(b). For the time per iteration, Butterfly-ADF is about $2\times$ slower than Butterfly-ADAM as it requires two gradient-like computation per iteration (See Line 7 and 8 of Algorithm 6). Surprisingly, we note that Butterfly-ALS and Butterfly-ADAM almost overlap with each other. This is likely due to the fact that the butterfly rank $r=11$ is a small constant independent of $n$. As a result, we conclude that ALS is the best algorithm among the three for butterfly completion of highly oscillatory matrices due to its fastest convergence (see Fig. 4 and Fig. 5) and time per iteration. On the other hand, QTT-ADF is much faster than QTT-ALS due to relatively large QTT ranks. Therefore, we conclude that ADF is the best algorithm among the three for QTT completion for highly oscillatory matrices. Note that although QTT completion requires less $|\Omega|$ than butterfly completion (see Fig. 6(b)), the time per iteration at $L=12$ is $3349$s for QTT-ADF but only $373$s for Butterfly-ALS. This time difference will be even larger for larger $L$. In terms of asymptotic behaviors, $|\Omega|$ scales as $\mathcal{O}(n\log n)$ for both butterfly ($r$ stays as a small constant) and QTT completion ($r=\mathcal{O}(n^{0.5})$) algorithms, and super-linearly for low-rank completion. As a result, the butterfly completion algorithms empirically achieve the lowest total asymptotic computational complexity, $\mathcal{O}(n\log^{3}n)$, among all the completion algorithms.
Moreover, we study the effect of rank estimate and $|\Omega|$ on the test error for QTT-ADF and Butterfly-ALS (i.e., the best performing algorithms out of the three QTT completion algorithms and the tree butterfly completion algorithms) using an example of $n=4096$ ($L=10$). For QTT completion, the actual rank with tolerance $\epsilon=10^{-3}$ is 54 and we set the rank estimates to $r=54,112$. For butterfly completion, the actual rank with tolerance $\epsilon=10^{-3}$ is 11 and we set the rank estimates to $r=11,20$. The test errors with varying observation ratios $|\Omega|/n^{2}$ are plotted in Fig. 7(a). For most values of observation ratios, the use of rank estimate higher than necessary leads to over-parameterization and hence worse test errors. For this Green’s function example, QTT completion typically requires lower observation ratios than butterfly completion (e.g., QTT-ADF reaches an error of 0.013 at $|\Omega|/n^{2}=0.03$ while Butterfly-ALS reaches an error of 0.017 at $|\Omega|/n^{2}=0.1$), but struggles to achieve the same level of test errors as the ratio increases.
@@
Figure 6: (a) Computational time per iteration and (b) $|\Omega|$ for a Green’s function example with $c=4$, L=814L=8-14 using the low-rank matrix completion, butterfly completion (ALS, ADAM, ADF), and QTT completion (ALS, ADAM, ADF). Butterfly completion: $|\Omega|=6rn\log n$, $r=11$ (i.e., the actual butterfly rank for $\bm{T}$ with tolerance $\epsilon=10^{-3}$). Low-rank completion: $|\Omega|=10rn$ with $r=57,134,366$ (i.e., the actual rank for $\bm{T}$ with tolerance $\epsilon=10^{-3}$). Note that in theory low-rank representation requires $r=\mathcal{O}(n)$ for sufficiently large $n$, but these numerical tests exhibit $r=\mathcal{O}(\sqrt{n})$ up to $n=16384$. QTT completion: $|\Omega|=16r^{2}\log n$ with $r=30,54,96$ (i.e., the actual QTT rank for $\bm{T}$ with tolerance $\epsilon=10^{-3}$). The QTT ranks scale approximately as $r=\mathcal{O}(n^{0.5})$.
Figure 7: Test errors with varying observation counts $|\Omega|$ and rank estimates $r$ for butterfly completion with ALS and QTT completion with ADF. (a) A Green’s function example with $c=4$, $L=10$. The actual rank for $\bm{T}$ with tolerance $\epsilon=10^{-3}$ is 10 for butterfly decomposition and 54 for QTT decomposition. (b) A Radon transform example with $c=4$, $L=8$. The actual rank for $\bm{T}$ with tolerance $\epsilon=10^{-3}$ is 10 for butterfly decomposition and 380 for QTT decomposition. We remark that for Radon transforms, QTT exhibits full ranks and requires $|\Omega|=n^{2}$ for accurate recovery.
@@ 6.2 Inverse wave scattering
To generate the Green’s function matrix $\bm{T}$, we leverage the simulation package WAVEFORM [WAVEFORM] to simulate the responses $p(\mathbf{x}^{o}_{j})$, $j=1,2,\ldots,n$ for each excitation function $f({\mathbf{x}})=\delta(\mathbf{x}-\mathbf{x}^{s}_{i})$. The frequency is set to $\omega=500\pi$ and $n$ is set to 1024 (i.e., $c=4$, $L=8$). Once the matrix is fully generated, we assume only $|\Omega|$ entries of the matrix are observed and our goal is to fully recover $\bm{T}$. For the transmission problem, we empirically set $|\Omega|=27n\log n$, $r=9$ for the butterfly completion Algorithm 1 and $r=120$ for the low-rank completion. For the reflection problem, we empirically set $|\Omega|=48n\log n$, $r=12$ for the butterfly completion Algorithm 1 and $r=120$ for the low-rank completion. The reconstructed matrices are plotted in the last two columns of Fig. 8. Clearly, the butterfly completion algorithm achieves significantly better reconstruction quality than the low-rank completion algorithm.
@@ 6.3 Radon transform
First, we consider an example of $n=1024$ and compare the reconstruction quality of the butterfly completion algorithm (Algorithm 1) and low-rank matrix completion algorithm. The real part of the full matrix $\bm{T}$ and the observed submatrix $\bm{T}_{\Omega}$ are shown in Fig. 9(a) and (b), where the number of observed entries is empirically set to $|\Omega|=30n\log n=307200$. The low-rank matrix completion algorithm uses rank estimate $r=400$, max iterations $t_{max}=20$, and the reconstructed full matrix $\bm{X}$ is shown in Fig. 9(c). The butterfly completion algorithm (Algorithm 1) uses rank estimate $r=10$, max iterations $t_{max}=20$ and the initial guess with $R=10$. The reconstructed full matrix $\bm{X}$ is shown in Fig. 9(d). As one can see, the reconstruction quality of Algorithm 1 is significantly better than that of the low-rank completion algorithm, given the same set of observed entries. This is no surprise as the low-rank completion algorithm requires $r=\mathcal{O}(n)$ and $|\Omega|=\mathcal{O}(n^{2})$ for a good reconstruction.
@@
Next, we examine the convergence of the completion algorithms. We first consider an example of $n=4096$ ($L=10$) using the ALS-based (Algorithm 1) and gradient-based (Algorithm 3) and ADF-based (Algorithm 6) butterfly completion with rank $r=12$ (the actual butterfly rank is 10 with tolerance $\epsilon=10^{-3}$). Note that the QTT completion algorithm is not considered as the full QTT rank will yield the algorithm prohibitively expensive (see later discussion of Fig. 11 and Fig. 7(b) for more detail). The number of observed entries is empirically set to $|\Omega|=7rn\log n$ for training and $|\Omega_{\mathrm{test}}|=n\log n$ for testing. As the reference, we also tested the low-rank matrix completion with rank $r=100$ or $200$, but using the same $|\Omega|$ as butterfly completion. The convergence history for the training and test errors are shown in Fig. 10. We set max iterations $t_{max}=20$ for all the ALS-based and ADF-based algorithms and $t_{max}=60$ for the gradient-based algorithm. As expected, the low-rank matrix completion algorithms do not converge due to both insufficient numbers of $r$ and $|\Omega|$. In contrast, the butterfly completion algorithms converge with a small rank estimate $r=12$. Specially, the ALS-based butterfly algorithm converges in 16 iterations, and the ADF-based butterfly algorithm is slightly worse, while the gradient-based butterfly algorithm takes more than 60 iterations to converge.
@@
Figure 10: The convergence of a Radon transform example with $n=4096$, $c=4$, $L=10$ using butterfly completion (ALS, ADAM, ADF) with rank $r=12$ (the actual butterfly rank for $\bm{T}$ with tolerance $\epsilon=10^{-3}$ is 10), $|\Omega|=7rn\log n=3440640$ and $|\Omega_{\rm test}|=rn\log n=491520$. (a) Training error. (b) Test error. The results using low-rank matrix completion with rank $r=100,400$ and the same $|\Omega|$ as butterfly completion are also plotted. Note that the actual rank for $\bm{T}$ with tolerance $\epsilon=10^{-3}$ is $r=4096$ (i.e., full rank) for low-rank decomposition.
@@
Figure 11: (a) Computational time per iteration and (b) $|\Omega|$ for the Radon transform example with $c=4$, L=614L=6-14 using the low-rank matrix completion, butterfly completion (ALS, ADAM, ADF) and QTT completion (ADF). Butterfly completion: $|\Omega|=6n\log n$, $r=10$. Low-rank completion: $|\Omega|=n^{2}$ , $r=n$ (this requires all the entries as $\bm{T}$ is full-rank). QTT completion: $|\Omega|=n^{2}$ (this requires all the entries as the actual $T$-rank for $\bm{T}$ is $r=140,380,1140$ for $L=6,8,10$).
@@
Next, we investigate the computational complexity of the low-rank matrix, QTT and butterfly completion algorithms for the radon transform matrices. As $L\in\{6,\dots,14\}$, we set the number of observed entries to $|\Omega|=n^{2}$ for low-rank completion and QTT completion as they need all the entries to be observed to converge to sufficient accuracy, and $|\Omega|=6rn\log n$ for the butterfly algorithms. For QTT completion, we only test QTT-ADF as the actual QTT ranks with tolerance $\epsilon=10^{-3}$) for $L=6,8,10$ are $140,380,1140$, which are almost full-rank and scale as $\mathcal{O}(n)$. The rank estimate is $r=10$ for the butterfly algorithms, $r=n$ for the low-rank algorithm, and $r=140,380$ for the QTT completion algorithm. We plot the CPU time per iteration for each algorithm in Fig. 11(a) and the number of observed entries $|\Omega|$ in Fig. 11(b). Similar to the Green’s function example, the time per iteration for ALS-based and ADAM-based butterfly algorithms almost overlap with each other, and ADF-based butterfly algorithm is about $2\times$ slower. Clearly, the butterfly algorithms achieve the quasi-linear asymptotic complexity as $\mathcal{O}(n\log^{3}n)$. In stark contrast, low-rank and QTT completion algorithms scale at least as $\mathcal{O}(n^{4})$ due to the unfavorable $|\Omega|=n^{2}$ and $r=\mathcal{O}(n)$ scaling. Note that at $n=1024$, it requires 5 hours per iteration for QTT completion. Obviously, if all the entries of the matrix are available, one should use low-rank or QTT decomposition instead of low-rank or QTT completion.
Finally, we study the effect of rank estimate and $|\Omega|$ on the test error for QTT-ADF and Butterfly-ALS using an example of $n=1024$ ($L=8$). For QTT-ADF, the actual rank with tolerance $\epsilon=10^{-3}$ is 380 and we set the rank estimates to $r=380$. For Butterfly-ALS, the actual rank with tolerance $\epsilon=10^{-3}$ is 10 and we set the rank estimates to $r=10,20$. The test errors with varying observation ratios $|\Omega|/n^{2}$ are plotted in Fig. 7(b). For most values of observation ratios, the use of rank estimate higher than necessary leads to over-parameterization and hence worse test errors. It is clear that QTT-ADF requires almost all matrix entries and hence does not represent an effective completion algorithm for Radon transform matrices.